OpenStudy (anonymous):

If the arc on a particular circle has a length of 12 units, and the circumference of the circle is 36, which is the degree measure of the arc?

OpenStudy (anonymous):

The circumference of the circle is 36 and there are 360 degress so the ratio of circumference to degrees is:\[\frac{ 36 }{ 360 }\] The arc measure is 12 so the ratio of its length to degree measure equals that. Let's say the arc degree measure is x. The equation would be:\[\frac{ 36 }{ 360 }=\frac{ 12 }{ x }\] See how the degree measures are both at the bottom. Now we just have to solve for x So yo divide the numerator 36 by 3 to get 12 so also divide the denominator 360 by 3 so you get 120. So x=120. The ANSWER IS: 120 DEGREES